OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y=___x + ___

OpenStudy (susanna):

y-y=m(x-x) plug in the point and slope, and solve

OpenStudy (susanna):

y-4=2(x-3)

OpenStudy (susanna):

y-4=2x-6 (getting rid of the parentheses first.

OpenStudy (anonymous):

........would it be....2x-2?

OpenStudy (susanna):

y=2x-2 yes

OpenStudy (anonymous):

yay!